OpenStudy (anonymous):

Find the value of "a" for the following expression 2(7a+1b)=10(10a+3b)

OpenStudy (anonymous):

2(7a+1b)=10(10a+3b) 14a+2b=100a+30b 2b=86a+30b -28b=86a -14b=43a -14/43b=a a=-14/43b I believe that's it.

OpenStudy (anonymous):

Thanks a lot.

OpenStudy (anonymous):

No problem